IvaHida74:

A carpenter is putting a skylight in a roof. If the roof measures 6x + 9 by 4x + 3 and the skylight measures 4x + 6 by 3x + 6, what is the area of the remaining roof after the skylight is built. Put your answer in factored form.

jhonyy9:

like an idea 6x +9 = 3(2x +3) in this way the roof measures 3(2x +3) by 4x +3 the skylight measures 4x +6 = 2(2x +3) so will be 2(2x +3) by 3x +6 and now than you consider that the skylight was built what will be the area of remaining roof so write in this way \[\frac{ \frac{ 3(2x +3) }{ 4x +3 } }{ \frac{ 2(2x +3) }{ 3(x +2) } } = \frac{ 3(2x +3) }{ 4x +3 }×\frac{ 3(x +2) }{ 2(2x +3) }\] any idea now ?